9. Develop a profit-and-loss statement for the Westgate division of North Industries. This division manufactures light fixtures sold to consumers through home improvement and hardware stores. Cost of goods sold represents 40% of sales. Marketing expenses include selling expenses, promotion expenses, and freight. Selling expenses include sales salaries totaling $3 million per year and sales commissions (5% of sales). The company spent $3 million on advertising last year, and freight costs were 10% of sales. Other costs include $2 million for managerial salaries and expenses for the marketing function and another $3 million for indirect overhead allocated to the division.

Develop the profit-and-loss statement. The sales revenue was $40 million last year from sale of 200,000 fixtures. Calculate Westgate’s breakeven sales.
Answer #1
PROFIT AND LOSS STATEMENT :
Sales $ 40,000,000.00
Less: Variable cost
Cost of goods sold (20*40%) $ 16,000,000.00
Sales commission (20*5%) $    2,000,000.00
Freight cost (20*10%) $    4,000,000.00 $ 22,000,000.00
Contribution margin $ 18,000,000.00
Less: Fixed expenses
Selling expenses $    3,000,000.00
Advertising expenses $    3,000,000.00
Managerial salaries $    2,000,000.00
Indirect overhead $    3,000,000.00 $ 11,000,000.00
Net Profit $   7,000,000.00
Break even sales = Fixed cost/contribution margin = 40,000,000/45% = $88,888,888.89
contribution margin ratio =18,000,000 /40,000,000 = 45%